An ideal gas has molar heat capacity Cv at constant volume. The gas undergo the process T=To(1-aV) where a, To is constant and V is volume. Then molar heat capacity of the gas is 

options are:
(a) Cv + R(1-aV)/aV
(B) Cv - R(1-aV)/aV
(c) Cv - R(1+aV)/2aV
(d) Cv + R(1+aV)/2aV

Grade:12

2 Answers

TANISHA
8 years ago
why aint i gettin a reply?
keshav
7 years ago
C=Cv+(P*dV) /(ndT) C=Cv-R[R-(1-$V)]/[$V]$= alpha Hope it helps!! Feel free to ask any question... We will help you out as much as possible!!